# Targeted Data Fusion for Causal Survival Analysis Under Distribution Shift

### v1

@@ 5 Discussion
In this work, we proposed two methods for estimating treatment-specific survival functions in a target population, leveraging external source data with possibly shifted covariate and outcome distributions. When multi-source data can be pooled and CCOD holds, our CCOD estimator is consistent, semiparametrically efficient, and handles covariate shifts. If CCOD is violated, the federated estimator remains consistent, offers moderate efficiency gains, and preserves data privacy by avoiding individual-level data sharing. This framework is also applicable to settings such as subgroup analysis (Yang et al., 2023) and multi-center clinical trials (Zhuang et al., 2024).
We note several limitations of our work. First, although our simulations show efficiency gains for the federated method, its optimality under data-sharing constraints is not established; exploring efficient covariate-adaptive weighting (Li & Luedtke, 2023) may be a promising future avenue. Second, when data sharing is allowed but CCOD may fail, it is unclear if any method outperforms both target-site-only and federated estimators, as our semiparametric efficiency relies on CCOD. Third, our time-specific federated weights, while flexible, have two drawbacks: (i) they may produce nonsmooth weight estimates over time, and (ii) they are computationally intensive in continuous time when working on a fine time grid. Therefore, future work should aim to more efficiently capture temporal trends and appropriately smooth these weights.
Our framework can potentially be extended to various settings, including surrogate markers (Gao et al., 2024a; Han et al., 2022), conformal inference (Liu et al., ), dynamic treatment regimes (Zhang et al., 2013; Chu et al., 2023), and data-driven individual selection from external sources (Zhu et al., 2024; Gao et al., 2024b). It may also apply to direct treatment effect modeling for survival outcomes, such as marginal structural (Yang et al., 2018) and structural nested models (Yang et al., 2020), as well as other survival estimands like restricted mean survival times (Han, 2023; Hua et al., 2024) and truncation contexts (Morenz et al., 2024; Han, 2024; Wang et al., 2024).
@@ Appendix A Additional Details and Results for Numerical Experiments
In this appendix, we present additional details and results for our numerical experiments in Section 3. All experiments in our study were performed using the statistical language R (version 4.4.2).
@@ A.1 Data generating process
with $\rho=1.2$, $\lambda=0.6$, and $U_{1},U_{2}\sim\text{Uniform}(0,1)$. This technique follows Austin (2012). Thus, the observed times and event indicators are $Y=\min(T,C),\Delta=\mathbb{I}(T\leq C)$, respectively. Under this data generating process (DGP), the event time is generated as a similar way of days in a year (365 days), and we truncate the censoring time at $t=200$ days to mimic the end of follow-up in survival analysis. Our DGP supports the following scenarios based on site-specific heterogeneity:
@@ Appendix B Additional Data Analysis Results
Figure 9: Estimated region-specific survival curves of the HVTN 704/HPTN 085 and HVTN 703/HPTN 081 trials. SA (our target region in the main text) and OA exhibit relatively similar curves, indicating less heterogeneity of these two regions. In contrast, both BP and US regions show significant differences to SA, which also confirms why they often have small or zero federated weights in Panel (C) of Figure 3. The BP and US also show a substantial difference on their curves.
Furthermore, in Figure 10, we present the results—including survival curve estimations and federated weights—using three regions other than SA as the target population. For the federated weights, similar to Figure 3 in the main text, we applied locally weighted regression (Cleveland & Devlin, 1988) to smooth the observed weights over the study period, providing a clearer visualization of temporal trends in this specific example.
@@
Appendix C Double/debiased Machine Learning Algorithm for Nuisance Function Predictions
In this section, we detail the procedures for fitting and predicting nuisance functions. Algorithm 2 outlines the DML algorithm employed in Algorithm 1 for federated causal survival estimation. A similar process is applied to the CCOD and other estimators when predicting their respective nuisance functions.
@@ D.1 Theory of the CCOD estimator
Recall the estimand $\theta^{0}(t,a)=\mathbb{E}\{S^{0}(t\mid a,\mathbf{X})\mid R=0\}=\mathbb{E}\{\bar{S}(t\mid a,\mathbf{X})\mid R=0\}$ under Assumption 2.6. We write $\bar{\Lambda}(t\mid a,\mathbf{x})=\int_{0}^{t}\frac{\bar{N}_{1}(du\mid a,\mathbf{x})}{\bar{D}(u\mid a,\mathbf{x})}$ where $\bar{N}_{\delta}(du\mid a,\mathbf{x})=\mathbb{P}(Y\leq t,\Delta=\delta\mid A=a,\mathbf{X}=\mathbf{x})$ and $\bar{D}(t\mid a,\mathbf{x})=\mathbb{P}(Y\geq t\mid A=a,\mathbf{X}=\mathbf{x})$. Then, by semiparametric theory, we write the equation
00 0 $=\frac{\partial}{\partial\epsilon}\theta^{0}(t,a)\bigg{|}_{\epsilon=0}=\frac{\partial}{\partial\epsilon}\mathbb{E}\{\bar{S}_{\epsilon}(t\mid a,\mathbf{X})\mid R=0\}\bigg{|}_{\epsilon=0}$
@@
with
$\frac{\partial}{\partial\epsilon}\bar{\Lambda}_{\epsilon}(du\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$ $=\dfrac{\frac{\partial}{\partial\epsilon}\bar{N}_{1,\epsilon}(du\mid a,\mathbf{x})\mid_{\epsilon=0}}{\bar{D}(u\mid a,\mathbf{x})}-\dfrac{\frac{\partial}{\partial\epsilon}\bar{R}_{\epsilon}(u\mid a,\mathbf{x})\mid_{\epsilon=0}\bar{N}_{1,\epsilon}(du\mid a,\mathbf{x})}{\bar{D}(u\mid a,\mathbf{x})^{2}}.$
@@
$\frac{\partial}{\partial\epsilon}\bar{N}_{1,\epsilon}(du\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$ $=\frac{\partial}{\partial\epsilon}\mathbb{P}_{\epsilon}(Y\leq u,\Delta=1\mid A=a,\mathbf{X}=\mathbf{x})\bigg{|}_{\epsilon=0}=\frac{\partial}{\partial\epsilon}\iint\mathbb{I}(y\leq u,\delta=1)\mathbb{P}_{\epsilon}(dy,d\delta\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$
$=\iint\mathbb{I}(y\leq u,\delta=1)\dot{\ell}(y,\delta\mid a,\mathbf{x},k)\mathbb{P}(dy,d\delta\mid a,\mathbf{x})=\int_{\delta}\mathbb{I}(\delta=1)\dot{\ell}(u,\delta\mid a,\mathbf{x},k)\mathbb{P}(du,d\delta\mid a,\mathbf{x}),$
@@
$\frac{\partial}{\partial\epsilon}\bar{D}_{\epsilon}(u\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$ $=\frac{\partial}{\partial\epsilon}\mathbb{P}_{\epsilon}(Y\geq u\mid A=a,\mathbf{X}=\mathbf{x})\bigg{|}_{\epsilon=0}=\frac{\partial}{\partial\epsilon}\iint\mathbb{I}(y\geq u)\mathbb{P}_{\epsilon}(dy,d\delta\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$
@@
$\quad+\iiiint\mathbb{I}(u\leq t,u\leq y)\frac{\bar{S}(t\mid a,\mathbf{x})\bar{S}(u-\mid a,\mathbf{x})}{\bar{S}(u\mid a,\mathbf{x})\bar{D}(u\mid\mathbf{x})}\dot{\ell}(y,\delta\mid a,\mathbf{x})\mathbb{P}(dy,d\delta\mid a,\mathbf{x})\bar{N}_{1}(du\mid a,\mathbf{x})\mu(d\mathbf{x})$
@@
$\quad+\iiint\bar{S}(t\mid a,\mathbf{x})\int_{0}^{t\wedge y}\frac{\bar{S}(u-\mid a,\mathbf{x})}{\bar{S}(u\mid a,\mathbf{x})\bar{D}(u\mid\mathbf{x})^{2}}\bar{N}_{1}(du\mid a,\mathbf{x})\dot{\ell}(y,\delta\mid a,\mathbf{x})\mathbb{P}(dy,d\delta\mid a,\mathbf{x})\mu(d\mathbf{x})$
$=\mathbb{E}\left[\bar{S}_{A}(t\mid\mathbf{X})\frac{I(A=a)}{\bar{\pi}(a\mid\mathbf{X})}\left\{\bar{H}(t\wedge Y,A,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)\bar{S}(Y-\mid A,\mathbf{X})}{\bar{S}(Y\mid A,\mathbf{X})\bar{D}(Y\mid A,\mathbf{X})}\right\}\dot{\ell}(Y,\Delta\mid A,\mathbf{X})\right],$
@@
$\mathbb{E}\{\bar{H}(t\wedge Y,A,\mathbf{X})\mid A=a,\mathbf{X}=\mathbf{x}\}$ $=\iint^{t}\mathbb{I}(u\leq y)\frac{\bar{S}(u-\mid a,\mathbf{x})\bar{N}_{1}(du\mid a,\mathbf{x})}{\bar{S}(u\mid a,\mathbf{x})\bar{D}(u\mid a,\mathbf{x})^{2}}\mathbb{P}(dy\mid a,\mathbf{x})$
@@
$\mathbb{E}\left[\bar{H}(t\wedge Y,A,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)\bar{S}(Y-\mid A,\mathbf{X})}{\bar{S}(Y\mid A,\mathbf{X})\bar{D}(Y\mid A,\mathbf{X})}~{}\bigg{|}~{}A=a,\mathbf{X}=\mathbf{x}\right]=0$
@@
$\frac{\partial}{\partial\epsilon}\theta^{0}(t,a)\bigg{|}_{\epsilon=0}$ $=\mathbb{E}\bigg{[}\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}[\bar{S}(t\mid a,\mathbf{X})-\theta^{0}(t,a)]\dot{\ell}(\mathcal{O})-\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\mathbb{E}\bigg{\{}\bar{S}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}$
@@
Therefore, an uncentered EIF of $\theta^{0}(t,a)$ at $\mathbb{P}$ is found as
$\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\bar{S},\bar{G},\bar{\pi})=\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P})$ $=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\bar{S}(t\mid a,\mathbf{X})-\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}\bar{S}(t\mid a,\mathbf{X})$
@@
D.1.2 Lemmata for the CCOD estimator

### v2

@@ 5 Discussion
We proposed two methods for estimating treatment-specific survival functions in a target population, leveraging external source data with possibly shifted covariate and outcome distributions. When multi-source data can be pooled and CCOD holds, our CCOD estimator is consistent, semiparametrically efficient, and handles covariate shifts. If CCOD is violated, the federated estimator remains consistent, offers moderate efficiency gains, and preserves patient privacy.
Limitations of our work present opportunities for future work. First, although our simulations show efficiency gains for the federated method, exploring efficient covariate-adaptive weighting [34] may be a promising future avenue. Second, when data sharing is allowed but CCOD may fail, it is unclear if any method outperforms both target-site-only and federated estimators, as our semiparametric efficiency is based on the CCOD assumption. Third, our time-specific federated weights, while flexible, may produce nonsmooth weights over time, and can be computationally intensive in continuous time. Therefore, future work should aim to more efficiently capture temporal trends and appropriately smooth these weights. Finally, future work should aim to incorporate time-varying covariates, which can improve efficiency by leveraging additional predictive information [40]. To the best of our knowledge, existing methods are limited to settings with discrete survival times [1, 19, 13], and unifying the related methodology for continuous time remains an open problem.
Our framework can potentially be extended to various settings, including surrogate markers [20, 28], conformal inference [35], dynamic treatment regimes [55, 9], and data-driven individual selection from external sources [56, 21]. It may also apply to direct treatment effect modeling for survival outcomes, such as marginal structural models [54] and structural nested models [53], as well as other survival estimands like restricted mean survival times [23, 30], competing risk [36] and truncation [37, 24, 47].
@@ A.1 Data generating process
with $\rho=1.2$, $\lambda=0.6$, and $U_{1},U_{2}\sim\text{Uniform}(0,1)$. This technique follows [2]. Thus, the observed times and event indicators are $Y=\min(T,C),\Delta=\mathbb{I}(T\leq C)$, respectively. Under this data generating process (DGP), the event time is generated as a similar way of days in a year (365 days), and we truncate the censoring time at $t=200$ days to mimic the end of follow-up in survival analysis. Our DGP supports the following scenarios based on site-specific heterogeneity:
@@ Appendix B Additional Data Analysis Results
Table 1 presents summary statistics for baseline covariates and outcomes in the AMP trial data, stratified by region and treatment group. Comparing the treatment groups—both overall and within each region—we observe that the treated group consistently shows a lower average event proportion. Additionally, some covariates appear to shift across regions; for example, among treated participants, the standardized risk scores exhibit notably different means when comparing SA to BP and US.
Treated (bnAb) group
Total ($n=3,076$) SA ($n=679$) OA ($n=608$) BP ($n=846$) US ($n=943$)
Age (year) at baseline 25.9 (4.60) 27.0 (5.19) 25.4 (4.59) 25.1 (3.70) 26.2 (4.68)
Standardized risk score 0.0 (1.00) -0.01 (1.00) 0.02 (1.00) 0.76 (0.67) -0.68 (0.71)
Weight at baseline (kg) 72.8 (15.64) 68.8 (14.24) 65.2 (12.63) 70.9 (12.42) 82.3 (16.43)
HIV diagnosis by week-80 107 (3.48%) 27 (3.98%) 20 (3.29%) 46 (5.44%) 14 (1.49%)
Control (placebo) group
Total ($n=1,535$) SA ($n=340$) OA ($n=297$) BP ($n=428$) US ($n=470$)
Age (year) at baseline 25.9 (4.72) 26.6 (5.28) 25.4 (4.78) 25.2 (3.94) 26.1 (3.79)
Standardized risk score 0.0 (1.00) 0.02 (0.92) -0.02 (0.98) 0.75 (0.67) -0.68 (0.73)
Weight at baseline (kg) 72.5 (16.35) 67.6 (14.77) 65.1 (13.64) 71.1 (12.84) 81.8 (17.5)
HIV diagnosis by week-80 67 (4.36%) 16 (4.71%) 13 (4.38%) 29 (6.78%) 9 (1.91%)
Table 1: Summary statistics of AMP trial data by treatment group and region. The standardized risk score is a baseline score built by machine learning models [12] that is predictive to the time-to-event outcome. Age, standardized risk score and weight are summarized by mean (standard deviation), while the HIV diagnosis by week-80 is summarized by count (proportion).
@@
Figure 10: Estimated region-specific survival curves of the HVTN 704/HPTN 085 and HVTN 703/HPTN 081 trials. SA (our target region in the main text) and OA exhibit relatively similar curves, indicating less heterogeneity of these two regions. In contrast, both BP and US regions show significant differences to SA, which also confirms why they often have small or zero federated weights in Panel (B) of Figure 4 in the main text. The BP and US also show a substantial difference on their curves.
Furthermore, in Figure 11, we present the results—including survival curve estimations and federated weights—using three regions other than SA as the target population. For the federated weights, similar to Figure 4 in the main text, we applied locally weighted regression [10] to smooth the observed weights over the study period, providing a clearer visualization of temporal trends in this specific example.
@@
Appendix C Details of the Proposed Algorithms
In this section, we describe (i) the procedure for federated estimation in Algorithm 1, and (ii) the DML procedure for fitting and predicting nuisance functions in Algorithm 2. Algorithm 1 implements the method introduced in Section 2.2 that is also illustrated in Figure 2. Algorithm 2 outlines the DML procedure used within Algorithm 1, and a similar approach is applied to the CCOD and other estimators when predicting their respective nuisance functions.
Algorithm 1 Federated Learning for Multi-Source Causal Survival Analysis.
1:  Input: Multi-source right-censored data $\{\mathcal{O}_{i}=(\mathbf{X}_{i},A_{i},Y_{i},\Delta_{i},R_{i}),i=1,\dots,n\}$, a time horizon $\tau>0$, and a fine time grid $\{0,\epsilon,2\epsilon,\dots,\tau\}$ for $[0,\tau]$, $\epsilon>0$.
2:  Output: Estimated treatment-specific survival curves $\widehat{\theta}^{\text{fed}}_{n}(t,a)$ and variance $\widehat{\mathcal{V}}_{t,a}=\widehat{\mathcal{V}}(\widehat{\theta}^{\text{fed}}_{n}(t,a))$ for $a\in\{0,1\}$ and $t\in\{0,\epsilon,2\epsilon,\dots,\tau\}$.
3:  for $(t,a)\in\{0,\epsilon,2\epsilon,\dots,\tau\}\times\{0,1\}$ do
8:  end for
9:  Return:
$\widehat{\theta}^{\text{fed}}_{n}(t,a)=\sum_{k=0}^{K-1}\widehat{\eta}^{k}_{t,a}\widehat{\theta}^{k,0}_{n}(t,a),\text{ and }\widehat{\mathcal{V}}_{t,a}^{\dagger}\text{ for all }(t,a)\in\{0,\epsilon,2\epsilon,\dots,\tau\}\times\{0,1\}.$
$\dagger$: $\widehat{\mathcal{V}}_{t,a}$ is specified in (D.3) in Appendix D.3 by replacing the fixed ${\boldsymbol{\eta}}_{t,a}$ with the estimated value $\widehat{\boldsymbol{\eta}}_{t,a}$.
@@ D.1 Theory of the CCOD estimator
Recall the estimand $\theta^{0}(t,a)=\mathbb{E}\{S^{0}(t\mid a,\mathbf{X})\mid R=0\}=\mathbb{E}\{\bar{S}(t\mid a,\mathbf{X})\mid R=0\}$ that follows from Assumption 2.5. We write $\bar{\Lambda}(t\mid a,\mathbf{x})=\int_{0}^{t}\frac{\bar{N}_{1}(du\mid a,\mathbf{x})}{\bar{D}(u\mid a,\mathbf{x})}$ where $\bar{N}_{\delta}(du\mid a,\mathbf{x})=\mathbb{P}(Y\leq t,\Delta=\delta\mid A=a,\mathbf{X}=\mathbf{x})$ and $\bar{D}(t\mid a,\mathbf{x})=\mathbb{P}(Y\geq t\mid A=a,\mathbf{X}=\mathbf{x})$.
Recall that a mean zero, finite variance function $\varphi^{*0}_{t,a}(\mathcal{O};\mathbb{P})$ is called an influence function of the target estimand (a functional) $\theta^{0}(t,a)=\theta^{0}(t,a;\mathbb{P})$ at $\mathbb{P}$ if, for any one-dimensional regular parametric submodel $\{\mathbb{P}_{\epsilon}:\epsilon\in[0,1)\}$ through $\mathbb{P}\equiv\mathbb{P}_{0}$,
$\left.\frac{\partial}{\partial\epsilon}\theta^{0}(t,a;\mathbb{P}_{\epsilon})\right|_{\epsilon=0}=\mathbb{E}_{\mathbb{P}}[\varphi^{*0}_{t,a}(\mathcal{O};\mathbb{P})\dot{\ell}(\mathcal{O})],$
where ˙$\dot{\ell}(\mathcal{O})$ is the score function of the submodel at $\epsilon=0$ (i.e., typically, ˙$\dot{\ell}(\mathcal{O})={\partial}\log{\{p_{\epsilon}(\mathcal{O})\}}/{\partial\epsilon}~{}|_{\epsilon=0}$), where $p_{\epsilon}(\cdot)$ denotes the probability density (likelihood) function under submodel $\mathbb{P}_{\epsilon}$ [4]. Thus, to find the EIF, we begin by writing the equation
00 0 $=\frac{\partial}{\partial\epsilon}\theta^{0}(t,a;\mathbb{P}_{\epsilon})\bigg{|}_{\epsilon=0}=\frac{\partial}{\partial\epsilon}\mathbb{E}_{\mathbb{P}_{\epsilon}}\{\bar{S}_{\epsilon}(t\mid a,\mathbf{X})\mid R=0\}\bigg{|}_{\epsilon=0}$
@@
where $\mu(\cdot)$ denotes the distribution of $\mathbf{X}$ induced by $\mathbb{P}$ and, for any sets of variables $V$ and $W$, ˙$\dot{\ell}_{V\mid W}$ denotes the conditional score function of $V$ given $W$, i.e., typically $\partial\log{\{p_{\epsilon}(V\mid W)\}}/{\partial\epsilon}~{}|_{\epsilon=0}$—note that such scores always satisfy $\mathbb{E}_{\mathbb{P}}(\dot{\ell}_{V\mid W}\mid W)=0$ [4].
@@
$\alpha\mapsto\bar{S}(t;H)\int_{0}^{t}\frac{\bar{S}(u-;H)}{\bar{S}(u;H)}\alpha(du)$
at $H$.
Thus, by letting $H(t)=\bar{\Lambda}_{\epsilon}(t\mid a,\mathbf{x})$ and the chain rule, the integrand in the second term becomes
@@
Furthermore,
$\frac{\partial}{\partial\epsilon}\bar{\Lambda}_{\epsilon}(du\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$ $=\dfrac{\frac{\partial}{\partial\epsilon}\bar{N}_{1,\epsilon}(du\mid a,\mathbf{x})\mid_{\epsilon=0}}{\bar{D}(u\mid a,\mathbf{x})}-\dfrac{\frac{\partial}{\partial\epsilon}\bar{D}_{\epsilon}(u\mid a,\mathbf{x})\mid_{\epsilon=0}\bar{N}_{1,\epsilon}(du\mid a,\mathbf{x})}{\bar{D}(u\mid a,\mathbf{x})^{2}}.$
@@
$\frac{\partial}{\partial\epsilon}\bar{N}_{1,\epsilon}(du\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$ $=\frac{\partial}{\partial\epsilon}\mathbb{P}_{\epsilon}(Y\leq u,\Delta=1\mid A=a,\mathbf{X}=\mathbf{x})\bigg{|}_{\epsilon=0}$
$=\frac{\partial}{\partial\epsilon}\iint\mathbb{I}(y\leq u,\delta=1)\mathbb{P}_{\epsilon}(dy,d\delta\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$
$=\iint\mathbb{I}(y\leq u,\delta=1)\dot{\ell}(y,\delta\mid a,\mathbf{x},k)\mathbb{P}(dy,d\delta\mid a,\mathbf{x})$
$=\int_{\delta}\mathbb{I}(\delta=1)\dot{\ell}(u,\delta\mid a,\mathbf{x},k)\mathbb{P}(du,d\delta\mid a,\mathbf{x}),$
@@
$\frac{\partial}{\partial\epsilon}\bar{D}_{\epsilon}(u\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$ $=\frac{\partial}{\partial\epsilon}\mathbb{P}_{\epsilon}(Y\geq u\mid A=a,\mathbf{X}=\mathbf{x})\bigg{|}_{\epsilon=0}$
$=\frac{\partial}{\partial\epsilon}\iint\mathbb{I}(y\geq u)\mathbb{P}_{\epsilon}(dy,d\delta\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$
@@
$\quad+\iiiint\mathbb{I}(u\leq t,u\leq y)\frac{\bar{S}(t\mid a,\mathbf{x})\bar{S}(u-\mid a,\mathbf{x})}{\bar{S}(u\mid a,\mathbf{x})\bar{D}(u\mid\mathbf{x})}$
×˙$\quad\qquad\times\dot{\ell}(y,\delta\mid a,\mathbf{x})\mathbb{P}(dy,d\delta\mid a,\mathbf{x})\bar{N}_{1}(du\mid a,\mathbf{x})\mu(d\mathbf{x})$
@@
$\quad+\iiint\bar{S}(t\mid a,\mathbf{x})\int_{0}^{t\wedge y}\frac{\bar{S}(u-\mid a,\mathbf{x})}{\bar{S}(u\mid a,\mathbf{x})\bar{D}(u\mid\mathbf{x})^{2}}\bar{N}_{1}(du\mid a,\mathbf{x})$
×˙$\quad\qquad\times\dot{\ell}(y,\delta\mid a,\mathbf{x})\mathbb{P}(dy,d\delta\mid a,\mathbf{x})\mu(d\mathbf{x})$
$=\mathbb{E}\left[\bar{S}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}\left\{\bar{H}(t\wedge Y,a,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)\bar{S}(Y-\mid a,\mathbf{X})}{\bar{S}(Y\mid a,\mathbf{X})\bar{D}(Y\mid a,\mathbf{X})}\right\}\dot{\ell}(Y,\Delta\mid a,\mathbf{X})\right],$
@@
$\mathbb{E}\{\bar{H}(t\wedge Y,A,\mathbf{X})\mid A=a,\mathbf{X}=\mathbf{x}\}$
$=\iint^{t}\mathbb{I}(u\leq y)\frac{\bar{S}(u-\mid a,\mathbf{x})\bar{N}_{1}(du\mid a,\mathbf{x})}{\bar{S}(u\mid a,\mathbf{x})\bar{D}(u\mid a,\mathbf{x})^{2}}\mathbb{P}(dy\mid a,\mathbf{x})$
@@
$\mathbb{E}\left[\bar{H}(t\wedge Y,A,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)\bar{S}(Y-\mid A,\mathbf{X})}{\bar{S}(Y\mid A,\mathbf{X})\bar{D}(Y\mid A,\mathbf{X})}~{}\bigg{|}~{}A,\mathbf{X}\right]=0$
almost surely. By the fact that score functions have (conditional) mean zero, and by the tower property of expectation, the above calculations implies that
@@
$\frac{\partial}{\partial\epsilon}\theta^{0}(t,a;\mathbb{P}_{\epsilon})\bigg{|}_{\epsilon=0}$
$=\mathbb{E}\bigg{[}\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}[\bar{S}(t\mid a,\mathbf{X})-\theta^{0}(t,a)]\dot{\ell}(\mathcal{O})-\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\mathbb{E}\bigg{\{}\bar{S}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}$
@@
Therefore, the (uncentered) EIF of the target estimand $\theta^{0}(t,a)$ at $\mathbb{P}$ is found as
$\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\bar{S},\bar{G},\bar{\pi})$ $=\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P})$
$=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\bar{S}(t\mid a,\mathbf{X})-\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}\bar{S}(t\mid a,\mathbf{X})$
@@
D.1.2 Regularity conditions for the CCOD estimator
The proof of Theorem 2.8 requires the following regularity conditions with $\bar{\pi}_{\infty}=\bar{\pi}$, $q^{0}_{\infty}=q^{0}$, $\bar{G}_{\infty}=\bar{G}$ and $\bar{S}_{\infty}=\bar{S}$.
Condition D.1.
There exists $\bar{\pi}_{\infty}$, $\bar{G}_{\infty}$, $\bar{\Lambda}_{\infty}$ and $\bar{S}_{\infty}$ such that
  1. (a)
    $\max_{m}\mathbb{P}\left[\frac{1}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}-\frac{1}{\bar{\pi}_{\infty}(a\mid\mathbf{X})}\right]^{2}\to_{p}0$,
  2. (b)
    $\max_{m}\mathbb{P}\left[{\widehat{q}^{0}_{m}(\mathbf{X})}-{q^{0}_{\infty}(\mathbf{X})}\right]^{2}\to_{p}0$,
  3. (c)
    $\max_{m}\mathbb{P}\left[\sup_{u\in[0,t]}\left|\frac{1}{\widehat{\bar{G}}_{m}(u\mid a,\mathbf{X})}-\frac{1}{\bar{G}_{\infty}(u\mid a,\mathbf{X})}\right|\right]^{2}\to_{p}0$, and
  4. (d)
Condition D.2.
There exists $\eta\in(0,\infty)$ such that for $\mathbb{P}$-almost all $\mathbf{x}$, $\widehat{\bar{\pi}}_{m}(a\mid\mathbf{x})\geq 1/\eta$, $\bar{\pi}_{\infty}(a\mid\mathbf{x})\geq 1/\eta$, $\widehat{q}^{0}_{m}(\mathbf{x})\geq 1/\eta$, $q^{0}_{\infty}(\mathbf{x})\geq 1/\eta$, $\widehat{\bar{G}}_{m}(t\mid a,\mathbf{x})\geq 1/\eta$, and $\bar{G}_{\infty}(t\mid a,\mathbf{x})\geq 1/\eta$ with probability tending to 1.
Condition D.3.
Define
$\bar{r}_{n,t,a,1}$ $=\max_{m}\mathbb{P}|\{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})-\bar{\pi}(a\mid\mathbf{X})\}\cdot\{\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})-\bar{S}(t\mid a,\mathbf{X})\}|,$
$\bar{r}_{n,t,a,2}$ $=\max_{m}\mathbb{P}|\{\widehat{q}^{0}_{m}(\mathbf{X})-q^{0}(\mathbf{X})\}\cdot\{\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})-\bar{S}(t\mid a,\mathbf{X})\}|,\text{ and}$
$\bar{r}_{n,t,a,3}$ $=\max_{m}\mathbb{P}\bigg{|}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\left\{\frac{\bar{G}(u\mid a,\mathbf{X})}{\widehat{\bar{G}}_{m}(u\mid a,\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\bigg{|}.$
Condition D.4.
Condition D.5.
D.1.3 Lemmata for the CCOD estimator